# Deep Learning-Assisted Weak Beam Identification in Dark-Field X-ray Microscopy

### v1

@@ Abstract
Dislocations control the mechanical behavior of crystalline materials, yet their quantitative characterization in bulk has remained elusive. Transmission Electron Microscopy provides atomic-scale resolution but is restricted to thin foils, limiting relevance to structural performance. Dark-field X-ray microscopy (DFXM) has recently opened access to three-dimensional, non-destructive imaging of dislocations in macroscopic crystals. A critical bottleneck, however, is the reliable identification of weak- versus strong-beam conditions. Weak-beam imaging enhances dislocation contrast, while strong-beam conditions are dominated by multiple scattering and obscure interpretation. Current practice depends on manual classification by specialists, which is subjective, slow, and incompatible with the scale of modern experiments. Here, we introduce a deep learning framework that automates this task using a lightweight convolutional neural network trained on small, hand-labeled datasets. By enabling robust, rapid, and scalable identification of imaging conditions, this approach transforms DFXM into a high-throughput tool, unlocking statistically significant studies of dislocation dynamics in bulk materials.
@@
Here we address a specific problem relevant for samples of high crystal quality, e.g. samples where individual dislocations within dislocation networks such as coherent small angle grain boundaries can be resolved. In such samples, strong multiple scattering occurs, and the dynamical theory of x-ray diffraction has to be applied in order to simulate the topographs observed near the maximum of the rocking curve [22, 23]. In the tails of the rocking curve, however, multiple scattering is negligible and the simpler kinematic theory of x-ray diffraction can be applied. This is called the WB condition. Therefore, the typical approach to analyzing RCI or DFXM data is to first segment the scanning data into weak- and strong-beam (SB) images, and them analyze and interpret the weak-beam images. The aim of this work is to provide a Machine Learning (ML) workflow for the automatic identification of weak-beam images in DFXM rocking curves. In subsequent data analysis steps, these weak beam images can the be used for the automatic identification of dislocations and other localized crystal defects.
@@ Results
To preserve all relevant information and enable reliable 3D reconstruction, we propose a patch-based approach for the localized identification of weak-beam conditions in each rocking layer. This method allows for small-patch-wise classification of beam conditions, a critical step for stacking and aligning the rocking layers. Ground truth labels for training were manually generated by sampling representative frames from the dataset and annotating them according to their corresponding beam condition.
For our the classification task of interest, we utilize a lightweight convolutional neural network (LCNN) based on a depthwise separable architecture [40, 41]. As depicted in Figure 2, the model operates on relatively small image patches pixels extracted from raw DFXM data.
@@
The data used in this study was derived from a previously published and publicly available DFXM dataset [13]. Each rocking step frame had a resolution of 2160×2560 pixels, which was divided into non-overlapping 64×64 patches. Labeling was performed manually at the image level, meaning that entire images were classified as either SB or WB conditions. Based on this, 6 WB images and 6 SB images were randomly selected, resulting in 6144 patches for each beam condition. The ground truth dataset was split into 80% for training and 20% for testing. All patches were log-scaled to enhance feature contrast and were passed through the neural network in mini-batches during training to ensure efficient learning and generalization.
The model is trained for 50 epochs, converging around epoch 30, as shown in Figure 3. Both training and testing losses decrease steadily, with no significant divergence, indicating effective generalization. The use of dropout further enhances model robustness by deactivating a subset of neurons during training epochs. Classification performance is evaluated using accuracy, defined as the proportion of correctly classified samples relative to the total dataset size:
where TP and TN denote true positives and true negatives, respectively, and FP and FN represent false positives and false negatives. Additionally, the final epoch confusion matrix provides additional insight, showing 272 correctly classified SB patches and 67 correctly classified WB patches, with only 28 SB and 1 WB instances misclassified.
To benchmark our LCNN and to see how much we gained in terms of accuracy versus the computation cost, we compare with ResNet18 [43] and VGG16 [44] the other lightweight variants of the Resnet and VGG families. Table 1 summarizes their respective performance in terms of parameter count, training time, and accuracy.
| Model | Trainable Parameters | Training Time (S) | Accuracy (%) | |||
| LCNN | 879 | K | 23 | .87 | 92 | .70 |
| ResNet18 | 11 | M | 75 | .07 | 97 | .13 |
| VGG16 | 134 | M | 208 | .09 | 89 | .10 |
@@ Discussion
After training, we proceed by assigning our LCNN to label the entire rocking curves by identifying WB regions, which are particularly of interest for dislocation mapping. To enhance defect visibility, frames corresponding to WB-labeled patches are extracted and combined into an integrated image. This approach excludes SB contributions dominated by dynamical diffraction. Figure 4 shows the integrated patches under both strong and weak beam conditions. Postprocessing is then applied to improve image contrast:
@@
Figure 4 (A) and (B) shows LCNN-labeled diffraction contrast maps under WB and SB conditions after post-processing. WB conditions enhance dislocation contrast, making dislocation boundaries and branching structures more distinct. During inference, each frame within a rocking curve is classified as WB or SB, allowing fast and consistent defect identification while excluding SB frames for efficient data reduction. Importantly, Figure 4 (C) and (D) demonstrate model predictions on previously unseen experimental data, where the LCNN generalizes successfully and correctly distinguishes WB from SB. The WB-labeled rocking layers are then assembled to reconstruct three-dimensional dislocation networks for automated and spatially resolved analysis of defect structures.
To place this work in context, it is instructive to compare our LCNN-based workflow with other relevant approaches for weak-beam condition identification in DFXM. For instance, Huang et al [25] proposed a method similar to principal component analysis (PCA) that uses Gram-Schmidt orthogonalization instead to decompose rocking curve data into three orthogonal components corresponding to the SB condition and the WB conditions on either side of the rocking curve. Unlike the data-driven approach employed here, their method does not rely on any training data but instead defines physically meaningful basis functions from manually selected regions of interest. The main advantage of such a physics-based orthogonalization is its interpretability and its grounding in the underlying diffraction contrast mechanisms. However, this also makes it more dependent on careful manual input and less straightforward to implement in an automated or real-time setting, particularly for large, heterogeneous datasets. By contrast, the lightweight convolutional neural network presented in this work enables fully automated, patch-level classification of beam conditions with minimal computational cost, allowing high-throughput processing of large volumes of DFXM data. The trade-off is that the LCNN requires representative training data and may need to be retrained when applied to new material systems with significantly different features.
@@
As a final validation of our approach, we reproduced the three-dimensional dislocation structures previously published in Ref. [13], using only the two-dimensional image layers automatically identified as WB conditions by our model. Figure 5 presents a direct comparison: panel (A) shows the reconstruction obtained through our fully automated workflow, while panel (B) corresponds to the manually segmented result from the original study. Despite using approximately ninety percent less data and requiring no human interpretation, the automated method yields excellent agreement with the published result, clearly demonstrating its reliability and efficiency. This capability is particularly relevant for time-resolved experiments at advanced synchrotron and free-electron laser sources, such as ESRF-EBS and XFELs, where large volumes of data are acquired rapidly.
To better understand the distinction between weak and strong beam conditions, we evaluated our model across multiple datasets. In certain cases, the separation between the two conditions is not clearly defined, making threshold-based methods unreliable. This emphasizes the advantage of our data-driven approach, which captures subtle, context-specific features beyond raw intensity. Our model offers an efficient and scalable solution for dislocation imaging across diverse materials and experimental configurations.
Conclustion and Outlook
@@
In this work, we presented an automated and computationally efficient framework for analyzing DFXM rocking curves using lightweight convolutional neural networks. Our model enables accurate and scalable identification of weak-beam conditions crucial for interpreting diffraction contrast and visualizing dislocations in crystalline materials. This approach significantly accelerates the analysis process, opening the door to high-throughput, three-dimensional dislocation mapping in bulk samples.
@@
Looking ahead, we aim to extend this method for real-time, on-the-fly data reduction during DFXM experiments, thereby streamlining acquisition and enabling adaptive imaging strategies. Concurrently, we are developing models to infer dislocation types and Burgers vectors by training neural networks on synthetic datasets generated from geometrical optics simulations [11, 46]. These advancements will contribute toward a more quantitative, interpretable, and fully automated framework for defect characterization in crystalline materials. The methodology presented here is also broadly applicable to other imaging modalities where weak-beam contrast plays a critical role, including X-ray topography, white-beam Laue microscopy, and transmission electron microscopy.
@@
The data used in this work originates from DFXM experiments conducted on three different samples. These experiments were performed at the beamlines ID06-HXM [52] and ID03 [15] of the European Synchrotron Radiation Facility (ESRF). The sample used in this work was a single crystal of industrially-pure aluminum (99.99%), purchased from the Surface Preparation Laboratory, with dimensions , oriented with the long axis perpendicular to the scattering plane. Prior to the experiment, the sample was annealed for 10 h at 590 ∘C in atmospheric air, then slowly cooled in the furnace. For further details see Ref. 13.

### v2

@@ Abstract
Dislocations control the mechanical behavior of crystalline materials, yet their quantitative characterization in bulk has remained elusive. Transmission Electron Microscopy provides atomic-scale resolution but is restricted to thin foils, limiting relevance to structural performance. Dark-field X-ray microscopy (DFXM) has recently opened access to three-dimensional, non-destructive imaging of dislocations in macroscopic crystals. A critical bottleneck, however, is the reliable identification of weak- versus strong-beam conditions. Weak-beam imaging enhances dislocation contrast, while strong-beam conditions are dominated by multiple scattering and obscure interpretation. Current practice depends on manual classification by specialists, which is subjective, slow, and incompatible with the scale of modern experiments. Here, we introduce a deep learning framework that automates this task using a lightweight convolutional neural network trained on small, hand-labeled datasets. By enabling robust, rapid, and scalable identification of imaging conditions, this approach supports scalable DFXM analysis, unlocking statistically significant studies of dislocation dynamics in bulk materials.
@@
Here we address a specific problem relevant for samples of high crystal quality, e.g. samples where individual dislocations within dislocation networks such as coherent small angle grain boundaries can be resolved. In such samples, strong multiple scattering occurs, and the dynamical theory of x-ray diffraction has to be applied in order to simulate the topographs observed near the maximum of the rocking curve [21, 22]. In the tails of the rocking curve, however, multiple scattering is negligible and the simpler kinematic theory of x-ray diffraction can be applied. This is called the WB condition. Therefore, the typical approach to analyzing rocking curve imaging (RCI) or DFXM data is to first segment the scanning data into weak- and strong-beam (SB) images, and then analyze and interpret the weak-beam images. The aim of this work is to provide a Machine Learning (ML) workflow for the automatic identification of weak-beam images in DFXM rocking curves. In subsequent data analysis steps, these weak beam images can the be used for the automatic identification of dislocations and other localized crystal defects.
@@ Results
To preserve all relevant information and enable reliable 3D reconstruction, we propose a patch-based approach for the localized identification of weak-beam conditions in each rocking layer. This method allows for small-patch-wise classification of beam conditions. Ground truth labels for training were manually generated by sampling representative frames from the dataset and annotating them according to their corresponding beam condition.
The complete 3D dataset analyzed in this work (originally reported in Ref. 13) comprises 9311 frames, each containing approximately 4 million pixels. From this volume, six representative frames, 3 weak beam and 3 strong beam (0.064 of the total dataset) were selected for manual annotation. Labeling was performed on a per-patch basis rather than per full frame, and each patch was classified as either WB or SB, resulting in a binary classification task. Because local lattice curvature and strain may cause spatial variations in diffraction conditions within a single projection, particular care was taken during annotation. Only patches exhibiting clearly distinguishable WB or SB contrast were selected for training, while ambiguous or transitional regions were excluded to avoid label noise. The selected frames span representative diffraction conditions across the rocking curve, ensuring that the training data captures realistic contrast variability while maintaining physically well-defined labels.
For our classification task of interest, we utilize a lightweight convolutional neural network (LCNN) based on a depthwise separable architecture [39, 40]. As depicted in Figure 2, the model operates on relatively small image patches pixels extracted from raw DFXM data.
@@
The data used in this study originate from two previously published and publicly available DFXM datasets corresponding to two separate samples: Sample 1 reported in Ref. 13 and Sample 2 reported in Ref. 42. Detailed descriptions of the samples and experimental conditions are provided in the Data Acquisition section. The model was trained using the dataset from Sample 1. Each rocking step frame had a resolution of 2160×2560 pixels, which was divided into non-overlapping 64×64 patches. Labeling was performed manually at the image level, meaning that entire images were classified as either SB or WB conditions. Based on this, 3 WB images and 3 SB images were randomly selected, resulting in 6144 patches for each beam condition. The ground truth dataset was split into 80% for training and 20% for testing and 1300 validation patches. All patches were log-scaled to enhance feature contrast and were passed through the neural network in mini-batches during training to ensure efficient learning and generalization.
The model is trained for 50 epochs, converging around epoch 30, as shown in Figure 3. Both training and testing losses decrease steadily, with no significant divergence, indicating effective model generalization. The use of dropout further enhances model robustness by deactivating a subset of neurons during training epochs. Classification performance is evaluated using accuracy, defined as the proportion of correctly classified samples relative to the total dataset size:
where TP and TN denote true positives and true negatives, respectively, and FP and FN represent false positives and false negatives.Additionally, the final epoch confusion matrix provides further insight, showing that the majority of the validation set was correctly classified. However, we observe that 10 additional WB patches were misclassified. A likely explanation is related to intensity variations under the SB condition. SB images often contain dynamical fringes and local intensity fluctuations. In regions where the WB contrast is weak or transitional (i.e., leaning toward SB), these variations can cause a WB patch to resemble SB at the patch level. This can bias the prediction toward SB and lead to the observed minor asymmetry. That said, the effect is subtle and mainly occurs in borderline regions rather than in clearly defined dislocation segments.
To benchmark our LCNN and evaluate the gains in accuracy relative to computational cost, we compared it with ResNet18 [43] and VGG16 [44], which are lightweight variants of the ResNet and VGG families. Table 1 summarizes their respective performance in terms of parameter count, training time, accuracy, and deployment time for labeling 300 rocking layers. The models were implemented using PyTorch, a Python deep learning library, and trained for 50 epochs with batches of 32 patches at a time on NVIDIA Tesla V100-SXM2-32GB GPUs.
| Model | Trainable Parameters | Training Time (S) | Accuracy (%) | Deployment Time (h) | ||||
| LCNN | 879 | K | 23 | .87 | 92 | .70 | 0 | .41 |
| ResNet18 | 11 | M | 75 | .07 | 97 | .13 | 1 | .32 |
| VGG16 | 134 | M | 208 | .09 | 89 | .10 | 1 | .55 |
@@
Although the field of view and pixel size differ between the 10 and 2 objectives, the weak-beam contrast associated with individual dislocations remains comparable in spatial extent, as illustrated in Fig. 4(a,b) for 10 and Fig. 4(c,d) for 2. Therefore, the selected 64×64 patch size does not require systematic rescaling with magnification. Only in cases of significantly higher dislocation density, where multiple strain fields overlap within a single patch, would a smaller patch size be advantageous.
@@ Discussion
After training, we proceed by assigning our LCNN to label the entire rocking curves by identifying WB regions, which are particularly of interest for dislocation mapping. To enhance defect visibility, frames corresponding to WB-labeled patches are extracted and combined into an integrated image. This approach excludes SB contributions dominated by dynamical diffraction. Figure 4 shows the integrated patches under both strong and weak beam conditions. Here, integrated WB and SB images are constructed by summing all rocking-curve patches identified as WB or SB, respectively, thereby producing a single representative WB image and SB image for further quantitative analysis and 3D reconstruction. Postprocessing is then applied to improve image contrast in the integrated WB images:
@@
Figure 4 (A) and (B) shows LCNN-labeled diffraction contrast maps under WB and SB conditions after post-processing. WB conditions enhance dislocation contrast, making dislocation boundaries and branching structures more distinct. Furthermore, Fig. 4 (C) and (D) demonstrate model predictions on previously unseen experimental data, where the LCNN correctly distinguishes WB from SB. Finally, the WB-labeled rocking layers are then assembled to reconstruct three-dimensional dislocation networks for spatially resolved analysis of defect structures.
To place this work in context, it is instructive to compare our LCNN-based workflow with other relevant approaches for weak-beam condition identification in DFXM. For instance, Huang et al [24] proposed a method similar to principal component analysis (PCA) that uses Gram-Schmidt orthogonalization (GSO) instead to decompose rocking curve data into three orthogonal components corresponding to the SB condition and the WB conditions on either side of the rocking curve. Unlike the data-driven approach employed here, their method does not rely on any training data but instead defines physically meaningful basis functions from manually selected regions of interest. The main advantage of such a physics-based orthogonalization is its interpretability and its grounding in the underlying diffraction contrast mechanisms. However, this also makes it more dependent on careful manual input and less straightforward to implement in an automated or real-time setting, particularly for large, heterogeneous datasets.
In the GSO formulation, the image stack is represented as a linear combination of three orthogonal basis states corresponding to WB-left, WB-right, and SB conditions. However, this orthogonality assumption is not generally satisfied in experimental datasets. In practice, WB-left and WB-right conditions often exhibit asymmetric contrast variations that depend on the local microstructure and diffraction condition, which can lead to incomplete separation and residual SB contamination in the extracted WB components. Furthermore, both PCA and GSO require manual selection of regions of interest to optimize the decomposition, typically repeated for individual 2D layers within a 3D scan. The outcome therefore becomes sample and operator dependent, and the procedure is time-intensive. In contrast, the LCNN approach is trained using image patches extracted from randomly selected frames within a dataset and can then be applied to the entire volume without further manual intervention. This ensures consistent treatment across the dataset and significantly reduces user bias. When reproducing the GSO analysis reported by Huang et al. [24], we observe that the GSO-derived WB component remains partially contaminated by residual SB fringes, whereas the LCNN output isolates dislocation contrast with substantially reduced SB background contributions. The LLCN presented in this work therefore enables automated, patch-level classification of beam conditions with minimal computational cost, supporting high-throughput processing of large DFXM volumes. The trade-off is that the LCNN requires representative training data and may need retraining when applied to material systems with substantially different contrast characteristics.
@@
As a final validation of our approach, we reproduced the three-dimensional dislocation structures previously published in Yildirim et al [13]. Figure 5 presents a direct comparison between the reconstruction reported by Yildirim et al., which is based on intensity thresholding applied to each individual 2D layer, and the reconstruction obtained using the LCNN approach proposed in this work. Fig. 5(A) shows the 3D reconstruction obtained through our fully automated workflow, whereas Fig. 5(B) corresponds to the manually segmented result from the original study after intensity thresholding. In Fig. 5(A), several individual dislocations are visible, which can be attributed to the integrated weak-beam approach that captures dislocation contrast under different weak-beam conditions throughout the rocking curve. At first glance, both approaches show similar features in the 3D volume in terms of dislocation boundaries. The visual similarity between the two volumes is further quantified in Fig. 5(C), which reports the distribution of symmetric nearest-neighbor distances between the reconstructed dislocation networks, with a mean error of approximately 2 m, indicating close spatial agreement. Fig. 5(D) displays representative dislocation boundaries extracted from both datasets, highlighting the consistent structural features recovered across multiple ROIs. Although the dislocation boundaries are reliably reconstructed, minor artifacts are also observed in the automated result, likely introduced by the patching method. Despite using approximately 90 less data and requiring no human intervention,The automated method shows good agreement of , corresponding to 5.4 million dislocation voxels compared to 4.2 million in the published results.This capability is particularly relevant for time-resolved experiments at advanced synchrotron and free-electron laser sources, such as ESRF-EBS and XFELs, where large volumes of data are acquired rapidly.
To better understand the distinction between WB and SB conditions, we evaluated our model across multiple datasets. While dislocation features are generally most evident under WB imaging conditions, lattice rotations introduced by dislocation boundaries frequently cause DFXM images to contain mixed WB and SB contributions within different regions of the same sample. As a result, in several rocking-curve positions, the intensity distributions of WB and SB frames partially overlap, particularly in transitional angular regions. This overlap complicates the interpretation of dislocation contrast and makes the separation between beam conditions ambiguous. Such ambiguity renders simple intensity threshold–based methods unreliable. Across the full dynamic range of detector intensities, variations in sensitivity to WB signal can introduce parasitic effects that further obscure the distinction between conditions. Consequently, manually selected WB frames may not be representative of the full scan range and can bias the reconstructed dislocation population.
This challenge is reflected quantitatively in the 3D reconstruction shown in Fig. 5. Within the same region of interest, conventional intensity thresholding yields approximately 4.2 million dislocation voxels (as reported in Ref. [13]), whereas the LCNN-based reconstruction identifies approximately 5.4 million voxels corresponding to an increase of about 25% in reconstructed dislocation volume. The improved completeness arises from integrating information across the full rocking curve rather than relying on a single global intensity criterion. These results show the advantage of our data-driven approach, which captures subtle, context-dependent features beyond raw intensity values and provides a scalable solution for dislocation imaging across diverse materials systems and experimental configurations.
Conclusion and Outlook
@@
In this work, we presented an automated and computationally efficient framework for analyzing DFXM rocking curves using lightweight convolutional neural networks. Our model enables accurate and scalable identification of weak-beam conditions crucial for interpreting diffraction contrast and visualizing dislocations in crystalline materials. This approach substantially reduces manual intervention and enables efficient three-dimensional dislocation mapping in bulk crystalline materials.
@@
Looking ahead, we aim to extend this method for real-time, on-the-fly data reduction during DFXM experiments, thereby streamlining acquisition and enabling adaptive imaging strategies. Concurrently, we are developing models to infer dislocation types and Burgers vectors by training neural networks on synthetic datasets generated from geometrical optics simulations [11, 45]. These advancements will contribute toward a more quantitative, interpretable, and fully automated framework for defect characterization in crystalline materials. The methodology presented here is also broadly applicable to other imaging modalities where WB contrast plays a critical role, including X-ray topography, white-beam Laue microscopy, and transmission electron microscopy.
@@
The data analyzed in this work originates from DFXM experiments conducted on two independent single-crystal aluminum samples during separate beam times at Beamline ID06-HXM (now ID03) of the European Synchrotron Radiation Facility (ESRF) [51, 15]. The first dataset corresponds to the study reported in Yildirim et al. [13], from which the full 3D dislocation reconstruction presented in Fig. 4(a-b) and Fig.5 is derived. We refer to this as Sample 1. In this case, the (002) Bragg reflection was mapped. The second dataset corresponds to the measurements reported in Borgi et al. [42], Fig. 4(c-d) where the (111) reflection was probed. We refer to this one as Sample 2.
The two datasets were collected under comparable optical configurations to assess both reconstruction performance and model transferability. Both samples were single crystals of industrially pure aluminum (99.99%), with dimensions , oriented with the long axis perpendicular to the scattering plane. Prior to the experiments, the samples were annealed for 10 h at 590 ∘C in atmospheric air and slowly cooled in the furnace to minimize residual strain. All DFXM measurements were performed using 17 keV photons selected by a Si(111) Bragg–Bragg double-crystal monochromator. The incident beam was vertically focused using a compound refractive lens (CRL) consisting of 58 one-dimensional Be lenslets with a radius of curvature m, yielding an effective focal length of 72 cm. The resulting beam profile at the sample was approximately line-focused, with a height of about 500 nm (FWHM) in the vertical direction and several hundred microns in the horizontal direction. This horizontal line beam illuminated a single plane within the crystal, thereby defining the microscope observation plane (Fig.1). Crystal alignment into the Bragg condition was performed using a near-field camera positioned 40 mm behind the sample. After alignment, the near-field detector was removed and the diffracted beam was magnified using an X-ray objective lens composed of 88 Be parabolic lenslets (two-dimensional focusing optics), each with m. The entry plane of the imaging CRL was positioned approximately 275-280 mm downstream of the sample along the diffracted beam path. The magnified image was projected onto a far-field detector positioned approximately 5 m from the sample. The far-field detector used an indirect detection scheme consisting of a scintillator, visible-light microscope optics, and a PCO.edge sCMOS camera. The visible optics could switch between 2 and 10 objectives. The present analysis focuses on the highest magnification configuration (10 visible objective), corresponding to an effective pixel size of approximately 40 nm along the laboratory direction and approximately 120 nm along the laboratory direction. Lower magnification (2) configurations correspond to proportionally larger effective pixel sizes.
In this work, we focused on rocking curve scans in both datasets. Rocking scans were acquired by rotating the tilt angle (see Fig. 1) about the laboratory axis over a defined angular range sampled in 30 steps. These one-dimensional scans of the rocking curve map components of the displacement gradient tensor field, including local strain and misorientation, enabling visualization of dislocations [13]. For the 3D reconstruction shown in Fig. 5 (Sample 1), rocking scans were collected for 301 -layers with 1 m spacing between successive slices. The rocking curve measurements were done over an angular range of with per step for Sample 1, totaling 31 images in the rocking curve. For sample 2, rocking curves were collected over a range of 0.02 with 0.002 per step. For both experiments, the exposure time was 1 seconds per frame. The resulting four-dimensional dataset was imported for analysis. For further details see Refs. 13, 42.